# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - Blockchain and Applications

### v2

@@ 22.1 Innovations and Challenges in Consensus Algorithms
Consensus mechanisms are the backbone of blockchain technology[380, 149]. They ensure that all participants in a decentralized network agree on the validity of transactions. Over the years, researchers have made significant advances in consensus algorithms to improve blockchain’s security, efficiency, and scalability. In this chapter, we explore various research developments and the challenges associated with these mechanisms.
@@ 22.1.1 In-Depth Research on Byzantine Fault Tolerance Algorithms
Byzantine Fault Tolerance (BFT) algorithms are designed to reach consensus in a distributed system[381], even when some participants behave maliciously or experience failures. This is essential for maintaining the integrity of decentralized systems like blockchains.
@@
One of the most prominent BFT algorithms is Practical Byzantine Fault Tolerance (PBFT)[382], introduced by Miguel Castro and Barbara Liskov in 1999. PBFT is efficient for smaller systems but becomes increasingly challenging to scale for larger networks, like public blockchains. Therefore, recent research has focused on optimizing BFT for blockchain scalability.
@@
For example, research on Tendermint[383], a popular BFT-based consensus mechanism, has shown promising results in improving scalability by reducing the communication overhead between nodes. In Tendermint, nodes communicate in rounds, and each round has a leader who proposes a block. If two-thirds of the nodes agree, the block is added to the blockchain. This reduces the time and complexity of reaching consensus, especially in networks with many participants.
@@ 22.1.2 Efficiency Optimization of Consensus Algorithms
Proof of Work (PoW)[380, 149] is the consensus algorithm used by Bitcoin. While PoW ensures security by requiring participants to solve complex mathematical puzzles, it consumes vast amounts of energy. Researchers are actively exploring ways to reduce this energy consumption. For instance, projects like Ethereum 2.0 are moving away from PoW and adopting Proof of Stake (PoS), which requires far less computational power.
One research direction involves reducing the complexity of the puzzles in PoW systems. By finding puzzles that still require work but are less energy-intensive, researchers aim to maintain security while lowering environmental impact[384].
@@
PoS is designed to replace the energy-hungry PoW by selecting validators based on the number of tokens they hold[231]. However, PoS can also suffer from scalability issues. To address this, researchers are investigating ways to improve PoS by reducing latency and increasing transaction throughput.
One approach is the use of sharding[232], where the blockchain is divided into smaller partitions, or ”shards.” Each shard processes transactions in parallel, which greatly increases the number of transactions the network can handle.
@@ 22.2.1 Variants and Optimizations of Practical Byzantine Fault Tolerance (PBFT)
Practical Byzantine Fault Tolerance (PBFT)[382] has proven effective for private or permissioned blockchain systems, where the number of nodes is limited and trust is partially established. However, scaling PBFT to public blockchains, which may have thousands of nodes, introduces challenges related to communication overhead and latency.
@@
Recent research has proposed several optimizations to make PBFT more scalable for large systems. For instance, in traditional PBFT, every node must communicate with every other node, which leads to quadratic communication costs[385]. Optimizations like reducing message complexity[386] have been proposed, where nodes only need to communicate with a subset of other nodes.
@@ 22.2.2 Research Progress in Zero-Knowledge Proof Consensus
Zero-knowledge proofs (ZKPs)[288] are cryptographic methods that allow one party to prove the validity of a statement without revealing the underlying data. These methods have gained significant attention in the context of blockchain because they can improve both privacy and scalability.
@@
One of the most exciting developments in ZKP is the concept of zero-knowledge rollups[387]. In this approach, multiple transactions are bundled together into a single proof, which is then verified on the main blockchain. This reduces the computational burden on the network, as fewer transactions need to be processed individually, thereby improving scalability without sacrificing security.
@@ 22.3 Research on Hybrid Consensus Algorithms
Hybrid consensus algorithms[286] combine multiple types of consensus mechanisms to leverage their respective strengths while mitigating their weaknesses. These algorithms are particularly useful in blockchain systems that require both high security and scalability.
@@ 22.3.1 Consensus Mechanisms Combining PoW and PoS
One promising area of research is the development of hybrid models that combine Proof of Work (PoW) and Proof of Stake (PoS)[388]. In these models, PoW is used to initially secure the network and prevent malicious attacks, while PoS is used to validate and approve transactions more efficiently.
@@
Decred[389] is an example of a cryptocurrency that uses a hybrid PoW/PoS consensus mechanism. In Decred, miners perform PoW to propose blocks, but stakeholders (those holding the coin) perform PoS to validate and approve these blocks. This combination ensures both the security benefits of PoW and the energy efficiency of PoS.
@@ 22.3.2 Research on Hybrid Consensus for Layer 1 and Layer 2
Layer 1 refers to the base layer of the blockchain, where fundamental operations like transaction processing and block creation take place. Layer 2 solutions are built on top of Layer 1 to improve scalability by offloading certain tasks[390].
Hybrid consensus mechanisms for Layer 1 and Layer 2 aim to combine the security of Layer 1 with the scalability of Layer 2. One popular solution is the Lightning Network[391], a Layer 2 technology for Bitcoin, where small transactions occur off-chain and are later settled on the main blockchain. Researchers are investigating how to combine Layer 1 and Layer 2 consensus in a way that maximizes both performance and security.
@@
Plasma[392] is another Layer 2 solution that uses a hybrid consensus approach. In Plasma, smart contracts and Merkle trees are used to create ”child chains” that run on top of the main blockchain. These child chains handle the majority of transactions, while the main chain is used for dispute resolution. This hybrid approach improves scalability by reducing the load on the main chain while ensuring that the child chains remain secure.
@@ 23.1.1 Comparison of On-Chain and Off-Chain Scaling
Blockchain scalability has been a major challenge due to the limitations in processing a large number of transactions while maintaining security and decentralization. Researchers have explored two primary methods for scaling blockchains: on-chain scaling and off-chain scaling[393].
@@
On-chain scaling involves making changes directly to the blockchain’s base layer[394]. A common approach is increasing the block size, which allows more transactions to be processed within each block. For instance, Bitcoin’s block size is capped at 1MB, which limits the number of transactions per second. By increasing this block size, we can allow more transactions to be processed at once.
@@
One well-known example is the Lightning Network for Bitcoin[395], which enables users to open payment channels and conduct numerous transactions without interacting directly with the Bitcoin blockchain. Off-chain solutions can significantly improve performance but often come with complexity in terms of security and usability.
@@
Both on-chain and off-chain scaling methods present trade-offs. On-chain scaling can compromise decentralization, as larger blocks may require more powerful nodes, leading to centralization[393]. Off-chain scaling, while maintaining decentralization, can introduce complexity and security risks, such as the need for participants to trust off-chain systems or mechanisms for dispute resolution[396].
@@ 23.1.2 The Theory and Implementation of Sharding
Sharding [232]is a scaling technique that aims to improve blockchain throughput by splitting the network into smaller pieces, called ”shards.” Each shard processes a subset of the total transactions in parallel, allowing multiple transactions to be processed simultaneously, thus increasing throughput.
@@
For example, in Ethereum’s sharding implementation[232], the network is divided into multiple shards, each functioning as a mini-blockchain. A central component known as the ”beacon chain” coordinates the activity of all shards to ensure security and integrity.
@@
Moreover, the communication between shards, known as ”cross-shard communication,”[397, 398] is a critical aspect of sharding. Ensuring that transactions between different shards are processed correctly is a key focus of current research. For instance, in Ethereum’s roadmap, sharding is planned to be implemented gradually to address these challenges and improve the network’s scalability.
@@ 23.2.1 Theoretical Foundations of State Channels
A state channel is established between two parties by locking some funds in a multi-signature wallet on the blockchain[399]. The participants then transact freely off-chain by exchanging signed messages that represent updates to the state of their agreement. When they decide to close the channel, they submit the final state of the channel to the blockchain for settlement.
@@ 23.2.2 Research Progress on Rollups
Rollups[243] are another Layer 2 solution gaining significant traction in blockchain research. Rollups bundle multiple transactions off-chain, compress them, and submit a summary to the main blockchain. There are two main types of rollups: Optimistic Rollups and Zero-Knowledge (ZK) Rollups.
@@
ZK-Rollups[245], on the other hand, generate cryptographic proofs (known as ZK-SNARKs) to verify the validity of every transaction. These proofs are submitted to the blockchain alongside the transaction summary, providing faster and more secure transaction verification. While ZK-Rollups provide more immediate security assurances, they require more complex computation to generate the cryptographic proofs.
@@ 23.3.1 Challenges of Cross-Chain Interoperability
Different blockchains often employ different consensus mechanisms, such as Proof of Work (PoW)[380, 149], Proof of Stake (PoS)[231], or others. These mechanisms can have different security assumptions and finality guarantees, making it difficult to synchronize state across chains.
@@ 23.3.2 Design of Decentralized Cross-Chain Bridges
One popular example is the ”Polkadot” network[265], which enables various blockchains to interoperate through a central relay chain. This allows different blockchains, known as ”parachains,”[400] to communicate with each other securely while maintaining their own consensus mechanisms.
@@ 23.4.1 Strategies for Improving Blockchain Throughput
For instance, Ethereum’s switch from Proof of Work to Proof of Stake in Ethereum 2.0[230] is expected to significantly increase transaction throughput by introducing more efficient block validation.
@@ 24.1.1 Academic Research on Zero-Knowledge Proofs
Zero-Knowledge Proofs (ZKPs)[288] are cryptographic protocols that enable one party (the prover) to convince another party (the verifier) that a statement is true without revealing any additional information about the statement. This concept, first introduced by Goldwasser, Micali, and Rackoff in the 1980s[401], has become essential in enhancing privacy in blockchain networks.
@@
Blockchains, by nature, are transparent. This is an excellent feature for security, but it raises privacy concerns since all transaction data is visible. ZKPs help solve this problem by allowing users to prove that they have the correct credentials or that they are authorized to perform an action without exposing sensitive details. For instance, in privacy-centric cryptocurrencies like Zcash[362], ZKPs (specifically zk-SNARKs, a variant of ZKPs[363]) are used to shield transaction amounts and addresses while still ensuring that the transaction is valid.
@@ 24.1.2 In-Depth Study of the MimbleWimble Protocol
MimbleWimble[402] is a blockchain protocol that enhances privacy and scalability by obfuscating transaction details and compressing the overall data. Introduced in 2016, MimbleWimble fundamentally restructures how transaction data is stored and validated.
@@ 24.1.3 Ring Signatures and Applications in Privacy Coins
Ring signatures[403] are cryptographic methods that allow a user to sign a message on behalf of a group, without revealing which specific member of the group signed the message. This technique is widely used in privacy-focused cryptocurrencies like Monero[366] to obfuscate transaction details.
@@
Monero[366] uses ring signatures to make transactions anonymous. When a user sends a transaction, their signature is grouped with several other signatures, creating a ”ring” of possible signers. As a result, it becomes impossible for an observer to tell which member of the ring made the transaction.
@@ 24.2.1 Applications of Multi-Party Computation (MPC) in Blockchain
Multi-Party Computation (MPC)[404] is a cryptographic protocol that allows multiple parties to jointly compute a function without revealing their private inputs. This ensures that even though computations are carried out, no party learns anything beyond the result of the computation.
@@
In blockchain, MPC can be used for secure voting[405], where participants can vote without revealing their choice, or for private transactions, where multiple parties can transact without revealing their individual transaction data.

### v3

@@ 22.1 Innovations and Challenges in Consensus Algorithms
Consensus mechanisms are the backbone of blockchain technology[gervais2016security, ferdous2020blockchainconsensusalgorithmssurvey]. They ensure that all participants in a decentralized network agree on the validity of transactions. Over the years, researchers have made significant advances in consensus algorithms to improve blockchain’s security, efficiency, and scalability. In this chapter, we explore various research developments and the challenges associated with these mechanisms.
@@ 22.1.1 In-Depth Research on Byzantine Fault Tolerance Algorithms
Byzantine Fault Tolerance (BFT) algorithms are designed to reach consensus in a distributed system[Zhang2022ReachingCI], even when some participants behave maliciously or experience failures. This is essential for maintaining the integrity of decentralized systems like blockchains.
@@
One of the most prominent BFT algorithms is Practical Byzantine Fault Tolerance (PBFT)[Ferenczi2022AnEV], introduced by Miguel Castro and Barbara Liskov in 1999. PBFT is efficient for smaller systems but becomes increasingly challenging to scale for larger networks, like public blockchains. Therefore, recent research has focused on optimizing BFT for blockchain scalability.
@@
For example, research on Tendermint[Buchman2018TheLG], a popular BFT-based consensus mechanism, has shown promising results in improving scalability by reducing the communication overhead between nodes. In Tendermint, nodes communicate in rounds, and each round has a leader who proposes a block. If two-thirds of the nodes agree, the block is added to the blockchain. This reduces the time and complexity of reaching consensus, especially in networks with many participants.
@@ 22.1.2 Efficiency Optimization of Consensus Algorithms
Proof of Work (PoW)[gervais2016security, ferdous2020blockchainconsensusalgorithmssurvey] is the consensus algorithm used by Bitcoin. While PoW ensures security by requiring participants to solve complex mathematical puzzles, it consumes vast amounts of energy. Researchers are actively exploring ways to reduce this energy consumption. For instance, projects like Ethereum 2.0 are moving away from PoW and adopting Proof of Stake (PoS), which requires far less computational power.
One research direction involves reducing the complexity of the puzzles in PoW systems. By finding puzzles that still require work but are less energy-intensive, researchers aim to maintain security while lowering environmental impact[Tsabary2019JustES].
@@
PoS is designed to replace the energy-hungry PoW by selecting validators based on the number of tokens they hold[Wendl2022TheEI]. However, PoS can also suffer from scalability issues. To address this, researchers are investigating ways to improve PoS by reducing latency and increasing transaction throughput.
One approach is the use of sharding[CortesGoicoechea2020ResourceAO], where the blockchain is divided into smaller partitions, or ”shards.” Each shard processes transactions in parallel, which greatly increases the number of transactions the network can handle.
@@ 22.2.1 Variants and Optimizations of Practical Byzantine Fault Tolerance (PBFT)
Practical Byzantine Fault Tolerance (PBFT)[Ferenczi2022AnEV] has proven effective for private or permissioned blockchain systems, where the number of nodes is limited and trust is partially established. However, scaling PBFT to public blockchains, which may have thousands of nodes, introduces challenges related to communication overhead and latency.
@@
Recent research has proposed several optimizations to make PBFT more scalable for large systems. For instance, in traditional PBFT, every node must communicate with every other node, which leads to quadratic communication costs[Yang2018LinBFTLB]. Optimizations like reducing message complexity[Bonomi2018MultihopBR] have been proposed, where nodes only need to communicate with a subset of other nodes.
@@ 22.2.2 Research Progress in Zero-Knowledge Proof Consensus
Zero-knowledge proofs (ZKPs)[ZHOU2024103678] are cryptographic methods that allow one party to prove the validity of a statement without revealing the underlying data. These methods have gained significant attention in the context of blockchain because they can improve both privacy and scalability.
@@
One of the most exciting developments in ZKP is the concept of zero-knowledge rollups[Capko2022StateOT]. In this approach, multiple transactions are bundled together into a single proof, which is then verified on the main blockchain. This reduces the computational burden on the network, as fewer transactions need to be processed individually, thereby improving scalability without sacrificing security.
@@ 22.3 Research on Hybrid Consensus Algorithms
Hybrid consensus algorithms[KumarJha2023HybridCM] combine multiple types of consensus mechanisms to leverage their respective strengths while mitigating their weaknesses. These algorithms are particularly useful in blockchain systems that require both high security and scalability.
@@ 22.3.1 Consensus Mechanisms Combining PoW and PoS
One promising area of research is the development of hybrid models that combine Proof of Work (PoW) and Proof of Stake (PoS)[Sun2019AnIH]. In these models, PoW is used to initially secure the network and prevent malicious attacks, while PoS is used to validate and approve transactions more efficiently.
@@
Decred[decredDecredMoney] is an example of a cryptocurrency that uses a hybrid PoW/PoS consensus mechanism. In Decred, miners perform PoW to propose blocks, but stakeholders (those holding the coin) perform PoS to validate and approve these blocks. This combination ensures both the security benefits of PoW and the energy efficiency of PoS.
@@ 22.3.2 Research on Hybrid Consensus for Layer 1 and Layer 2
Layer 1 refers to the base layer of the blockchain, where fundamental operations like transaction processing and block creation take place. Layer 2 solutions are built on top of Layer 1 to improve scalability by offloading certain tasks[Hafid2020ScalingBA].
Hybrid consensus mechanisms for Layer 1 and Layer 2 aim to combine the security of Layer 1 with the scalability of Layer 2. One popular solution is the Lightning Network[lightningLightningNetwork], a Layer 2 technology for Bitcoin, where small transactions occur off-chain and are later settled on the main blockchain. Researchers are investigating how to combine Layer 1 and Layer 2 consensus in a way that maximizes both performance and security.
@@
Plasma[ethereumPlasmaChains] is another Layer 2 solution that uses a hybrid consensus approach. In Plasma, smart contracts and Merkle trees are used to create ”child chains” that run on top of the main blockchain. These child chains handle the majority of transactions, while the main chain is used for dispute resolution. This hybrid approach improves scalability by reducing the load on the main chain while ensuring that the child chains remain secure.
@@ 23.1.1 Comparison of On-Chain and Off-Chain Scaling
Blockchain scalability has been a major challenge due to the limitations in processing a large number of transactions while maintaining security and decentralization. Researchers have explored two primary methods for scaling blockchains: on-chain scaling and off-chain scaling[Kim2018ASO].
@@
On-chain scaling involves making changes directly to the blockchain’s base layer[Pawar2020ASO]. A common approach is increasing the block size, which allows more transactions to be processed within each block. For instance, Bitcoin’s block size is capped at 1MB, which limits the number of transactions per second. By increasing this block size, we can allow more transactions to be processed at once.
@@
One well-known example is the Lightning Network for Bitcoin[Seres2019TopologicalAO], which enables users to open payment channels and conduct numerous transactions without interacting directly with the Bitcoin blockchain. Off-chain solutions can significantly improve performance but often come with complexity in terms of security and usability.
@@
Both on-chain and off-chain scaling methods present trade-offs. On-chain scaling can compromise decentralization, as larger blocks may require more powerful nodes, leading to centralization[Kim2018ASO]. Off-chain scaling, while maintaining decentralization, can introduce complexity and security risks, such as the need for participants to trust off-chain systems or mechanisms for dispute resolution[Dong2018CelerNB].
@@ 23.1.2 The Theory and Implementation of Sharding
Sharding [CortesGoicoechea2020ResourceAO]is a scaling technique that aims to improve blockchain throughput by splitting the network into smaller pieces, called ”shards.” Each shard processes a subset of the total transactions in parallel, allowing multiple transactions to be processed simultaneously, thus increasing throughput.
@@
For example, in Ethereum’s sharding implementation[CortesGoicoechea2020ResourceAO], the network is divided into multiple shards, each functioning as a mini-blockchain. A central component known as the ”beacon chain” coordinates the activity of all shards to ensure security and integrity.
@@
Moreover, the communication between shards, known as ”cross-shard communication,”[Das2020EfficientCT, Liu2024CHERUBIMAS] is a critical aspect of sharding. Ensuring that transactions between different shards are processed correctly is a key focus of current research. For instance, in Ethereum’s roadmap, sharding is planned to be implemented gradually to address these challenges and improve the network’s scalability.
@@ 23.2.1 Theoretical Foundations of State Channels
A state channel is established between two parties by locking some funds in a multi-signature wallet on the blockchain[Dziembowski2019MultipartyVS]. The participants then transact freely off-chain by exchanging signed messages that represent updates to the state of their agreement. When they decide to close the channel, they submit the final state of the channel to the blockchain for settlement.
@@ 23.2.2 Research Progress on Rollups
Rollups[Thibault2022BlockchainSU] are another Layer 2 solution gaining significant traction in blockchain research. Rollups bundle multiple transactions off-chain, compress them, and submit a summary to the main blockchain. There are two main types of rollups: Optimistic Rollups and Zero-Knowledge (ZK) Rollups.
@@
ZK-Rollups[Fernando2023PosterWA], on the other hand, generate cryptographic proofs (known as ZK-SNARKs) to verify the validity of every transaction. These proofs are submitted to the blockchain alongside the transaction summary, providing faster and more secure transaction verification. While ZK-Rollups provide more immediate security assurances, they require more complex computation to generate the cryptographic proofs.
@@ 23.3.1 Challenges of Cross-Chain Interoperability
Different blockchains often employ different consensus mechanisms, such as Proof of Work (PoW)[gervais2016security, ferdous2020blockchainconsensusalgorithmssurvey], Proof of Stake (PoS)[Wendl2022TheEI], or others. These mechanisms can have different security assumptions and finality guarantees, making it difficult to synchronize state across chains.
@@ 23.3.2 Design of Decentralized Cross-Chain Bridges
One popular example is the ”Polkadot” network[pokladotPolkadotInteroperable], which enables various blockchains to interoperate through a central relay chain. This allows different blockchains, known as ”parachains,”[Scott2023BringingTA] to communicate with each other securely while maintaining their own consensus mechanisms.
@@ 23.4.1 Strategies for Improving Blockchain Throughput
For instance, Ethereum’s switch from Proof of Work to Proof of Stake in Ethereum 2.0[Asif2023ShapingTF] is expected to significantly increase transaction throughput by introducing more efficient block validation.
@@ 24.1.1 Academic Research on Zero-Knowledge Proofs
Zero-Knowledge Proofs (ZKPs)[ZHOU2024103678] are cryptographic protocols that enable one party (the prover) to convince another party (the verifier) that a statement is true without revealing any additional information about the statement. This concept, first introduced by Goldwasser, Micali, and Rackoff in the 1980s[goldreich1994definitions], has become essential in enhancing privacy in blockchain networks.
@@
Blockchains, by nature, are transparent. This is an excellent feature for security, but it raises privacy concerns since all transaction data is visible. ZKPs help solve this problem by allowing users to prove that they have the correct credentials or that they are authorized to perform an action without exposing sensitive details. For instance, in privacy-centric cryptocurrencies like Zcash[zZcashPrivacyprotecting], ZKPs (specifically zk-SNARKs, a variant of ZKPs[Nitulescu2020zkSNARKsAG]) are used to shield transaction amounts and addresses while still ensuring that the transaction is valid.
@@ 24.1.2 In-Depth Study of the MimbleWimble Protocol
MimbleWimble[cointelegraphWhatMimblewimble] is a blockchain protocol that enhances privacy and scalability by obfuscating transaction details and compressing the overall data. Introduced in 2016, MimbleWimble fundamentally restructures how transaction data is stored and validated.
@@ 24.1.3 Ring Signatures and Applications in Privacy Coins
Ring signatures[Mercer2016PrivacyOT] are cryptographic methods that allow a user to sign a message on behalf of a group, without revealing which specific member of the group signed the message. This technique is widely used in privacy-focused cryptocurrencies like Monero[getmoneroMoneroProject] to obfuscate transaction details.
@@
Monero[getmoneroMoneroProject] uses ring signatures to make transactions anonymous. When a user sends a transaction, their signature is grouped with several other signatures, creating a ”ring” of possible signers. As a result, it becomes impossible for an observer to tell which member of the ring made the transaction.
@@ 24.2.1 Applications of Multi-Party Computation (MPC) in Blockchain
Multi-Party Computation (MPC)[Evans2019API] is a cryptographic protocol that allows multiple parties to jointly compute a function without revealing their private inputs. This ensures that even though computations are carried out, no party learns anything beyond the result of the computation.
@@
In blockchain, MPC can be used for secure voting[Sharma2019ScalableOS], where participants can vote without revealing their choice, or for private transactions, where multiple parties can transact without revealing their individual transaction data.